Windsor Inc. wishes to accumulate $1,240,000 by December 31, 2027, to retire bonds outstanding. The company deposits $200,000 on December 31, 2017, which will earn interest at 11% compounded quarterly, to help in the retirement of this debt. In addition, the company wants to know how much should be deposited at the end of each quarter for 10 years to ensure that $1,240,000 is available at the end of 2027. (The quarterly deposits will also earn at a rate of 11%, compounded quarterly.
Answer #1

Function Arguments PMT Rate 119%/4 = 0.0275 Nper 10*4 个=40 Pv 200000 - 一200000 个 = 1240000 number Fv 1240000 Type -9092.774937 Calculates the payment for a loan based on constant payments and a constant interest rate. Type is a logical value: payment at the beginning of the period 1; payment at the end of the period 0 or omitted. Formula result = S (9,092.77) Help on this function ок Cancel

Deposit at the end of each quarter is $9,092.77 or $9,093